OpenStudy (anonymous):

i need help with this question plz step by step (4+x)/3+4=(x-6)/2-6

OpenStudy (phi):

is it this? \[ \frac{4+x}{3} + 4 = \frac{x-6}{2}-6 \]

OpenStudy (anonymous):

yes

OpenStudy (phi):

The very first thing is add 6 to both sides of the equation. can you do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do you meen multiply by 6?

OpenStudy (phi):

write the equation on a piece of paper. on the left side write +6 . on the right side write +6 then combine 4+6 to get 10 combine -6+6 to get 0

OpenStudy (phi):

you should get \[ \frac{4+x}{3} + 10 = \frac{x-6}{2} \]

OpenStudy (phi):

now multiply both sides by 6. that means multiply every term by 6.

OpenStudy (anonymous):

\[\LARGE \frac{4+x}{3}+4+6=\frac{x-6}{2}\;\;\underbrace{-6+6}_{\text{This is 0}}\] this is what phi is telling you :)

OpenStudy (anonymous):

i get it ... go on

OpenStudy (phi):

now multiply both sides by 6. that means multiply every term by 6.

OpenStudy (anonymous):

now it is(4+x)+60=x-6 right?

OpenStudy (anonymous):

\[ \LARGE \frac{6(4+x)}{3}+6\cdot10=\frac{6(x-6)}{2}\]

OpenStudy (phi):

what happened to the 3 and the 2? They are still there in the bottom

OpenStudy (phi):

But notice you can divide 3 into 6 to get 2. and on the other side divide 2 into 6

OpenStudy (anonymous):

but didnt you get rid of that after multipling by 6?

OpenStudy (phi):

Only after you divide.

OpenStudy (anonymous):

ok.

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

4+x+60=x-6

OpenStudy (anonymous):

is that right?

OpenStudy (phi):

Do you see Kresh's post. Start with that. Divide the 3 into the 6. You should get 2

OpenStudy (anonymous):

\[\LARGE \frac{6(4+x)}{3}+6\cdot 10=\frac{6(x-6)}{2}\] transform it into... \[\LARGE \frac{6}{3}\cdot (4+x)+6\cdot 10=\frac{6}{2}\cdot (x-6)\]

OpenStudy (anonymous):

2*(4+x)+60=3*(x-6) does it look like that once simplified?

OpenStudy (phi):

looks good

OpenStudy (phi):

now we get rid of the parens. Distribute the 2. That is fancy way of saying, multiply 2 times each term in the parens.

OpenStudy (phi):

Do the same on the right hand side, distribute the 3

OpenStudy (anonymous):

like this 8+2x+60=3x-18

OpenStudy (phi):

yes. Now combine the 8 and 60

OpenStudy (anonymous):

2x+3x=-18-60-8 is that right?

OpenStudy (phi):

close. You did the numbers correctly. but to move the 3x to the other side you should add -3x to both sides.

OpenStudy (anonymous):

wait isnt it 2x-3x? the answer is 86

OpenStudy (anonymous):

that's correct.

OpenStudy (anonymous):

thank you both sooooooooooooooo much !!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

thank phi... he helped you, I just enjoyed watching lol :F

OpenStudy (anonymous):

thank you phi yr awesome !!!!!!!!!!!!!!!!!!!!!!

OpenStudy (phi):

So the part that confused you was "clearing the fraction" by multiplying by 6. You have to multiply by 6, but not ignore the denominators. You then divide the denominator into the 6. That makes them "go away"

OpenStudy (anonymous):

im not normally a dumb student but i needed help badly cuz my math teacher entred me in a math contest in may and i knew i couldnt go unltil i understood this:)

OpenStudy (phi):

May is coming up fast...

OpenStudy (anonymous):

thats the sad part i suck at hurrying up

OpenStudy (phi):

Good luck. But practice does help.

OpenStudy (anonymous):

is it possible in the future you could assist again i meanon this site?

OpenStudy (anonymous):

once again thanksssssssssssssssssssssssssss